How do I work this ??
QUESTION 3 Not yet answered Marked out of 3.00 PFlag question Using the Basic Accounting Equation Henderson Company had beginning-of-the-year total assets of $300,000 and total liabilities of $180,000. a. If during the year total assets increased by $15,000 and total liabilities increased by $40,000, what is the end-of-year total stockholders equity? $95,000 b. If during the year total assets increased by $60,000 and total liabilities decreased by $5,000, what is the end-of-year total stockholders equity? $ 185,000 c. If during the year total liabilities increased by $40,000 and total stockholders equity increased by $35,000, what are the end-of-year total assets?
Answer #1

Stockholders' equity = Total assets - Total liabilities

a) End of year total stockholders' equity = (300,000 + 15,000) - (180,000 + 40,000)

= $95,000

b) End of year total stockholders' equity = (300,000 + 60,000) - (180,000 - 5000)

= $185,000

c) Beginning of the year stockholders' equity = 300,000 - 180,000

= $120,000

Stockholders' equity = Total assets - Total liabilities

At the end of the year,

(120,000 + 35,000) = Total assets - (180,000 + 40,000)

155,000 = Total assets - 220,000

Total assets = 155,000 + 220,000

= $375,000